OpenStudy (amtran_bus):

what is 2 the squareroot of 5 squared?

OpenStudy (amtran_bus):



zepdrix (zepdrix):

Since both terms are included inside the brackets, you have to square both of them. \[\huge (2 \sqrt5)^2=2^2 \sqrt5^2=\qquad 2*2*\sqrt5*\sqrt5\]

OpenStudy (anonymous):